Grade 12th passPhysical Chemistry

1-What pressure difference (in kPa) must be generated across the length of a 20 cm vertical drinking straw in
order to drink a water like liquid of density 1.0 gm m–3. (g = 10 m/s2
2= A gas cylinder of 5 lit. capacity contains air under a pressure of 20 atm at 300 K. The air is to inflate cycle tubes
to a capacity of 3.5 lit. under a pressure of 4 atm. How many tubes can be inflated at the same temperature

vikas kumar
11 Years agoGrade 12th pass

1 Answer

Sunil Kumar FP
11 Years ago
1.pressure is given by=dgh,where
d is the density
g is the acceleration due to gravity,h is the height
density,d=10^-3*10*.2=2*10^-3 Pa


2)here the concept is that the number of mole of gas will remain same.let k be the number of cycle tube
number of mole in gas cylinder= k*number of mole in tube
5*20/R*300=k*3.5*4/300
k=7 tubes